# Causal Time Series Generation via Diffusion Models

### v1

@@ Abstract
Time series generation (TSG) synthesizes realistic sequences and has achieved remarkable success. Among TSG, conditional models generate sequences given observed covariates, however, such models learn observational correlations without considering unobserved confounding. In this work, we propose a causal perspective on conditional TSG and introduce causal time series generation as a new TSG task family, formalized within Pearl’s causal ladder, extending beyond observational generation to include interventional and counterfactual settings. To instantiate these tasks, we develop CaTSG, a unified diffusion-based framework with backdoor-adjusted guidance that causally steers sampling toward desired interventions and individual counterfactuals while preserving observational fidelity. Specifically, our method derives causal score functions via backdoor adjustment and the abduction–action–prediction procedure, thus enabling principled support for all three levels of TSG. Extensive experiments on both synthetic and real-world datasets show that CaTSG achieves superior fidelity and also supporting interventional and counterfactual generation that existing baselines cannot handle. Overall, we propose the causal TSG family and instantiate it with CaTSG, providing an initial proof-of-concept and opening a promising direction toward more reliable simulation under interventions and counterfactual generation.
@@ 1 Introduction
While introducing Int- and CF-TSG expands conditional TSG toward decision-oriented synthesis, it also raises several challenges: ground truth dilemma, invisible confounders, and evaluation difficulty. First, real-world time series often originate from human systems (e.g., mobility, health), where running interventions is ethically and practically infeasible (Xia et al., 2025a), thus causal supervision is rare and most datasets are observational (Matthay & Glymour, 2022). Second, confounders that simultaneously influence conditions and outcomes (e.g., individual difference and latent randomness) are typically unobserved and complex, making it non-trivial to model or adjust for them. Third, evaluating the quality of generated time series under these two tasks is difficult. Unlike images, time series lack reliable visual inspection, and their quality depends on preserving subtle statistical patterns (Narasimhan et al., 2024).
@@ 2 Background
Conditional Time Series Generation. Conditional TSG aims to synthesize realistic time series that conform to specified auxiliary information. Early methods like conditional GANs (Esteban et al., 2017; Smith & Smith, 2020) conditioned on discrete labels, but suffered from instability and mode collapse (Chen, 2021). Recent works adopt diffusion-based models (Sohl-Dickstein et al., 2015; Ho et al., 2020) with side information injected into the reverse process, enabling stable conditional generation (Tashiro et al., 2021). To handle heterogeneous metadata, TimeWeaver (Narasimhan et al., 2024) extends diffusion to categorical, continuous, and time-varying inputs. Other methods leverage structured state-space models (Alcaraz & Strodthoff, 2023), score-based transformers (Yuan & Qiao, 2024), or constrained objectives (Coletta et al., 2023). To improve adaptability across domains, recent efforts explore prompt-based control (Huang et al., 2025), disentanglement of dynamics and conditions (Bao et al., 2024), and textual guidance (Li et al., 2025). Consequently, diffusion models are currently a prevalent choice for time series generation (Yang et al., 2024; Narasimhan et al., 2024).
Diffusion Models & Score Function. Diffusion probabilistic models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song & Ermon, 2019) define a generative process by reversing a fixed forward noising process. Score function is one interpretation that diffusion models can also be interpreted as score-based generative models (Song & Ermon, 2019), which estimate the gradient of the log-density of noisy samples: $s_{t}(x_{t})=\nabla_{x_{t}}\log p(x_{t})$. Under the Gaussian forward process, the data score satisfies $\nabla_{x_{t}}\log q(x_{t}\mid x_{0})=-\frac{1}{\sigma_{t}}\,\epsilon,\quad\sigma_{t}=\sqrt{1-\bar{\alpha}_{t}}$. Accordingly, learning to predict the noise is equivalent to estimating the score up to a time-dependent scaling factor. In the conditional setting, the score becomes
@@
Causal Modeling for Generation. Causal views recently integrated into deep learning tasks e.g., representation learning (Yang et al., 2021), vision-language grounding (Wang et al., 2025), and anomaly detection (Lin et al., 2022; Xia et al., 2025b). In generative modeling (Komanduri et al., 2023), it has been explored across several modalities. In vision, models like CausalGAN (Kocaoglu et al., 2018) and Counterfactual Generative Networks (Sauer & Geiger, 2021) enable interventions and recombinations of disentangled factors for counterfactual image synthesis. In text generation, structural causal models guide interventions on latent narrative factors (Hu & Li, 2021), while in knowledge graphs, causal reasoning has been leveraged to generate hypothetical relations that enhance link prediction (Liu et al., 2021). Despite these advances, causal modeling for generation remains largely unexplored in time series.
@@ 3 A Causal View for Conditional Time Series Generation
Causal Ladder for TSG. Pearl’s Causal Ladder organizes reasoning into three levels: association, intervention, and counterfactual. We adapt this hierarchy to the context of TSG, summarized in Table 1. Level 1 (association) captures statistical correlations between observed contextual variables $C$ (e.g., temperature) and the outcome $X$ (e.g., traffic); Level 2 (intervention) models the outcome under controlled changes to $C$ while adjusting for potential confounding from unobserved common causes; and Level 3 (counterfactual) simulates alternate outcomes $X^{\prime}$ had $C$ been different, given the observed sample $X$ and $C$. This structured perspective enables us to unify and extend TSG tasks along the causal hierarchy.
General Problem Statement. Let $\mathcal{D}=\{(\mathbf{x},\mathbf{c})\}_{i=1}^{N}$ denote a dataset consisting of $N$ samples of multivariate time series $\mathbf{x}\in\mathbb{R}^{T\times D}$ and paired contextual sequences $\mathbf{c}\in\mathbb{R}^{T\times D_{c}}$. Each $\mathbf{x}=[\bm{x}_{1},\dots,\bm{x}_{T}]$ represents a length-$T$ time series over $D$ dimensions, and $\mathbf{c}$ contains auxiliary features in the same time span with dimension $D_{c}$. Our goal is to learn a generative model ${\mathcal{F}}_{\theta}$ such that the samples generated from ${\mathcal{F}}_{\theta}(\mathbf{c})$ match a target distribution of the form $P(\mathbf{x}^{\star}\mid{\color[rgb]{0.28125,0.4609375,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.28125,0.4609375,1}\mathcal{I}})$, where the form of the conditioning variable $\mathcal{I}$ determines the underlying generation task. Specifically, we consider:
@@ 4 Methodology
To instantiate $\mathcal{F}_{\theta}$ for the causal generative tasks, we first introduce a unified Structural Causal Model (SCM) (Pearl et al., 2000) as the theoretical foundation (Figure 2). We then derive the interventional and counterfactual objectives using two classical tools (Pearl et al., 2016): backdoor adjustment for Int-TSG and abduction–action–prediction procedure for CF-TSG. Given that diffusion models Ho et al. (2020) have become a powerful technique for time series modeling with strong performance and and stable training (Yang et al., 2024), we adopt diffusion as the backbone. Guided by the causal principles, within the diffusion framework, we derive a causal score function that guides the sampling process, and introduce CaTSG (Figure 3), a generative model that instantiates causal objectives in practice.
@@ 4.3 A Unified Causality-Guided Diffusion Framework
Framework Overview (Figure 3). CaTSG integrates three core components: (1) an environment inference module (EnvInfer $q_{\phi}$) that outputs latent representations $\mathbf{h}$, logits $\mathbf{s}$, and posterior environment weights $\mathbf{w}$ (the first two outputs used only for $\mathcal{L}_{\text{sw}}$), (2) a learnable environment bank $\mathbf{E}$ that provides representative embeddings of diverse environments, regularized by $\mathcal{L}_{\text{orth}}$, and (3) a denoising network (Denoiser $\varepsilon_{\theta}$) trained under a backdoor-adjusted formulation to regress Gaussian noise. Together, these modules instantiate the causal score function: EnvInfer assigns inputs $(\mathbf{x},\mathbf{c})$ to anchor embeddings in $\mathbf{E}$, and Denoiser predicts noise conditioned on $\mathbf{c}$ and $\mathbf{E}$, as well as unconditional predictions to follow a Classifier-Free Guidance (CFG)-style scheme (Ho & Salimans, 2022). We next detail the training and generation procedure, while the architectural specifications of EnvInfer and the Denoiser are provided in Appendix C.4.
@@ 5.5 Ablation Study
Figure 6: #Steps v.s. performance on Harmonic-VM.
@@ B.1 Causal Concepts
We now present the causal foundations relevant to our work. We begin by introducing Structural Causal Models (SCMs) (Pearl et al., 2000; Pearl, 2009), followed by a brief review of the rules of do-calculus (Pearl et al., 2016), which enable formal reasoning under interventions. We then derive the backdoor adjustment formula (Figure 2c) and introduce the three-step process of counterfactual generation (Figrue 2d).
@@
To compute such queries, taking our case as an example (Figure 2d), SCMs follow a principled three-step procedure (Pearl et al., 2000):
@@ B.2 Diffusion Models and Conditional Diffusion Models
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are a class of probabilistic generative models that learn to generate data through a two-phase procedure: a forward diffusion process that progressively adds noise to the data, followed by a reverse denoising process that learns to recover the original data distribution. In recent years, diffusion models have become a powerful paradigm for time series modeling, achieving strong performance across a wide range of tasks (Tashiro et al., 2021; Wen et al., 2023; Yuan & Qiao, 2024; Yang et al., 2024; Lin et al., 2024). We begin by introducing the unconditional setting, detailing both the forward and reverse processes. We then introduce the extension of this framework to the conditional setting, where external factors are incorporated to guide the generation process.
@@ B.2.1 Diffusion Models
Diffusion models can also be interpreted as a special class of score-based generative models (Song & Ermon, 2019), which aim to learn the gradient of the log-density of the data at each timestep. This gradient, known as the score function, indicates the direction in which the noisy input $x_{t}$ should be adjusted to increase its likelihood under the marginal distribution $p_{t}(x_{t})$. The score function is formally defined as:
@@ B.2.2 Conditional Diffusion Models
Ho & Salimans (2022) propose Classifier-Free Guidance (CFG), which can be interpreted as a reparameterization of classifier guidance by manipulating the gradient expression of the conditional log-probability. Starting from Bayes’ rule, we can write the second term in Eq 22 as:
@@ C.1 Motivation for a Diffusion Backbone
Our choice of a diffusion backbone is driven by the goals of Int-TSG and CF-TSG. First, under the score-based view, diffusion models learn and approximate scores $s(x,c)=\nabla_{x}\log p(x\mid c)$ (Eq. 1). Given identifiability of the interventional target $p(x\mid\mathrm{do}(c))$ instead of $p(x\mid c)$, our objective naturally extends to the interventional score $s(x,\mathrm{do}(c))=\nabla_{x}\log p(x\mid\mathrm{do}(c))$, which can be derived from the backdoor-adjusted formulation (Eq. 2). The derivation will be introduced in following sections. Second, diffusion models have become a powerful paradigm for time series modeling, achieving strong performance across imputation, generation, and forecasting tasks (Tashiro et al., 2021; Wen et al., 2023; Yuan & Qiao, 2024; Yang et al., 2024; Lin et al., 2024), offering stable training and good coverage compared to likelihood-free adversarial alternatives. These properties make diffusion a pragmatic and principled backbone for our causal TSG agenda.
@@ C.2 Derivation of backdoor-adjusted score function
The summation over latent confounder $e$ is inside the logarithm, preventing direct application of the gradient operator. This is a well-known difficulty in score-based inference under marginalization. To address this, we adopt a classical result often referred to as the log-derivative of a marginal or softmax trick (Mnih & Gregor, 2014), given by:
@@ D.1 Training objective & optimization
In addition, to encourage diversity among environment embeddings, we regularize the learnable environment bank $\mathbf{E}$ with an orthogonality loss:
@@ E.3 Empirical Evidence of Spurious Correlations
Real-world datasets. For the Air Quality dataset (Figure 12a and 12b), the conditional relation with temperature shows moderate between–split shifts, and wind direction exhibits pronounced nonmonotonic drift, both evident in the LOWESS curves and confirmed by the binned means. For the Traffic dataset (Figure 12c and 12d), the association with cloud coverage shows mild shifts, whereas the relation with hourly rainfall is more strongly nonstationary across splits. Note that the missing bins in Figure 12d arise because rain_1h is zero–inflated with a heavy–tailed positive part; under equal–width binning many mid–range intervals contain no observations and are omitted, producing the visible gap.
@@ H.6 Comparison with Related Works
Prior causal generative models are mostly developed for static modalities, e.g., images via label-graph–consistent GANs or modular recombinations (Kocaoglu et al., 2018; Sauer & Geiger, 2021), text via interventions on latent narrative factors (Hu & Li, 2021), and knowledge graphs via hypothetical triple generation (Liu et al., 2021). our work, to the best of our knowledge, is the first attempt to propose and instantiate causal time series generation, where causal queries target entire sequential data under interventions and regime shifts. Methodologically, within an SCM-based formulation, we realize causality through a back-door–adjusted score guidance that injects confounder adjustment directly into the diffusion sampling field, and we couple it with a learnable environment bank to model latent environments that drive nonstationarity. This is different from designs that either intervene on observed, single-shot factors or require acyclicity-constrained latent DAGs as in CausalVAE (Yang et al., 2021).

### v2

@@ Abstract
Time series generation (TSG) synthesizes realistic sequences and has achieved remarkable success. Among TSG, conditional models generate sequences given observed covariates, however, such models learn observational correlations without considering unobserved confounding. In this work, we propose a causal perspective on conditional TSG and introduce causal time series generation as a new TSG task family, formalized within Pearl’s causal ladder, extending beyond observational generation to include interventional and counterfactual settings. To instantiate these tasks, we develop CaTSG, a unified diffusion-based framework with backdoor-adjusted guidance that causally steers sampling toward desired interventions and individual counterfactuals while preserving observational fidelity. Specifically, our method derives causal score functions via backdoor adjustment and the abduction–action–prediction procedure, thus enabling principled support for all three levels of TSG. Extensive experiments on both synthetic and real-world datasets show that CaTSG achieves superior fidelity and also supporting interventional and counterfactual generation that existing baselines cannot handle. Overall, we propose the causal TSG family and instantiate it with CaTSG, providing an initial proof-of-concept and opening a promising direction toward more reliable simulation under interventions and counterfactual generation. Our source code is available at https://github.com/microsoft/TimeCraft/tree/main/CaTSG.
@@ 1 Introduction
While introducing Int- and CF-TSG expands conditional TSG toward decision-oriented synthesis, it also raises several challenges: ground truth dilemma, invisible confounders, and evaluation difficulty. First, real-world time series often originate from human systems (e.g., mobility, health), where running interventions is ethically and practically infeasible (Xia et al., 2025a), thus causal supervision is rare and most datasets are observational (Matthay and Glymour, 2022). Second, confounders that simultaneously influence conditions and outcomes (e.g., individual difference and latent randomness) are typically unobserved and complex, making it non-trivial to model or adjust for them. Third, evaluating the quality of generated time series under these two tasks is difficult. Unlike images, time series lack reliable visual inspection, and their quality depends on preserving subtle statistical patterns (Narasimhan et al., 2024).
@@ 2 Background
Conditional Time Series Generation. Conditional TSG aims to synthesize realistic time series that conform to specified auxiliary information. Early methods like conditional GANs (Esteban et al., 2017; Smith and Smith, 2020) conditioned on discrete labels, but suffered from instability and mode collapse (Chen, 2021). Recent works adopt diffusion-based models (Sohl-Dickstein et al., 2015; Ho et al., 2020) with side information injected into the reverse process, enabling stable conditional generation (Tashiro et al., 2021). To handle heterogeneous metadata, TimeWeaver (Narasimhan et al., 2024) extends diffusion to categorical, continuous, and time-varying inputs. Other methods leverage structured state-space models (Alcaraz and Strodthoff, 2023), score-based transformers (Yuan and Qiao, 2024), or constrained objectives (Coletta et al., 2023). To improve adaptability across domains, recent efforts explore prompt-based control (Huang et al., 2025), disentanglement of dynamics and conditions (Bao et al., 2024), and textual guidance (Li et al., 2025). Consequently, diffusion models are currently a prevalent choice for time series generation (Yang et al., 2024; Narasimhan et al., 2024).
Diffusion Models & Score Function. Diffusion probabilistic models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song and Ermon, 2019) define a generative process by reversing a fixed forward noising process. Score function is one interpretation that diffusion models can also be interpreted as score-based generative models (Song and Ermon, 2019), which estimate the gradient of the log-density of noisy samples: $s_{t}(x_{t})=\nabla_{x_{t}}\log p(x_{t})$. Under the Gaussian forward process, the data score satisfies $\nabla_{x_{t}}\log q(x_{t}\mid x_{0})=-\frac{1}{\sigma_{t}}\,\epsilon,\quad\sigma_{t}=\sqrt{1-\bar{\alpha}_{t}}$. Accordingly, learning to predict the noise is equivalent to estimating the score up to a time-dependent scaling factor. In the conditional setting, the score becomes
@@
Causal Modeling for Generation. Causal views recently integrated into deep learning tasks e.g., representation learning (Yang et al., 2021), vision-language grounding (Wang et al., 2025), and anomaly detection (Lin et al., 2022; Xia et al., 2025b). In generative modeling (Komanduri et al., 2023), it has been explored across several modalities. In vision, models like CausalGAN (Kocaoglu et al., 2018) and Counterfactual Generative Networks (Sauer and Geiger, 2021) enable interventions and recombinations of disentangled factors for counterfactual image synthesis. In text generation, structural causal models guide interventions on latent narrative factors (Hu and Li, 2021), while in knowledge graphs, causal reasoning has been leveraged to generate hypothetical relations that enhance link prediction (Liu et al., 2021). Despite these advances, causal modeling for generation remains largely unexplored in time series.
@@ 3 A Causal View for Conditional Time Series Generation
Causal Ladder for TSG. The causal ladder organizes reasoning into three levels: association, intervention, and counterfactual. We adapt this hierarchy to the context of TSG, summarized in Table 1. Level 1 (association) captures statistical correlations between observed contextual variables $C$ (e.g., temperature) and the outcome $X$ (e.g., traffic); Level 2 (intervention) models the outcome under controlled changes to $C$ while adjusting for potential confounding from unobserved common causes; and Level 3 (counterfactual) simulates alternate outcomes $X^{\prime}$ had $C$ been different, given the observed $X$ and $C$. This structured view enables us to unify and extend TSG tasks along the causal hierarchy.
General Problem Statement. Let $\mathcal{D}=\{(\mathbf{x},\mathbf{c})\}_{i=1}^{N}$ denote a dataset consisting of $N$ samples of multivariate time series $\mathbf{x}\in\mathbb{R}^{T\times D}$ and paired contextual sequences $\mathbf{c}\in\mathbb{R}^{T\times D_{c}}$. Each $\mathbf{x}=[\bm{x}_{1},\dots,\bm{x}_{T}]$ represents a length-$T$ time series over $D$ dimensions, and $\mathbf{c}$ contains auxiliary features in the same time span with dimension $D_{c}$. Our goal is to learn a generative model ${\mathcal{F}}_{\theta}$ such that the samples generated from ${\mathcal{F}}_{\theta}(\mathbf{c})$ match a target distribution of the form $P(\mathbf{x}^{\star}\mid{\color[rgb]{0.28125,0.4609375,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.28125,0.4609375,1}\mathcal{I}})$, where the form of the conditioning variable $\mathcal{I}$ determines the underlying generation task:
@@ 4 Methodology
To instantiate $\mathcal{F}_{\theta}$ for the causal generative tasks, we first introduce a unified Structural Causal Model (SCM) (Pearl and others, 2000) as the theoretical foundation (Figure 2). We then derive the interventional and counterfactual objectives using two classical tools (Pearl et al., 2016): backdoor adjustment for Int-TSG and abduction–action–prediction procedure for CF-TSG. Given that diffusion models Ho et al. (2020) have become a powerful technique for time series modeling with strong performance and and stable training (Yang et al., 2024), we adopt diffusion as the backbone. Guided by the causal principles, within the diffusion framework, we derive a causal score function that guides the sampling process, and introduce CaTSG (Figure 3), a generative model that instantiates causal objectives in practice.
@@ 4.3 A Unified Causality-Guided Diffusion Framework
Framework Overview (Figure 3). CaTSG integrates three core components: (1) an environment inference module (EnvInfer $q_{\phi}$) that outputs latent representations $\mathbf{h}$, logits $\mathbf{s}$, and posterior environment weights $\mathbf{w}$ (the first two outputs used only for $\mathcal{L}_{\text{sw}}$), (2) a learnable environment bank $\mathbf{E}$ that provides representative embeddings of diverse environments, regularized by $\mathcal{L}_{\text{orth}}$, and (3) a denoising network (Denoiser $\varepsilon_{\theta}$) trained under a backdoor-adjusted formulation to regress Gaussian noise. Together, these modules instantiate the causal score function: EnvInfer assigns inputs $(\mathbf{x},\mathbf{c})$ to anchor embeddings in $\mathbf{E}$, and Denoiser predicts noise conditioned on $\mathbf{c}$ and $\mathbf{E}$, as well as unconditional predictions to follow a Classifier-Free Guidance (CFG)-style scheme (Ho and Salimans, 2022). We next detail the training and generation procedure, while the architectural specifications of EnvInfer and the Denoiser are provided in Appendix C.4.
@@ 5.5 Ablation Study
Table 4: #Steps v.s. performance on Harmonic-VM.
@@ B.1 Causal Concepts
We now present the causal foundations relevant to our work. We begin by introducing Structural Causal Models (SCMs) (Pearl and others, 2000; Pearl, 2009), followed by a brief review of the rules of do-calculus (Pearl et al., 2016), which enable formal reasoning under interventions. We then derive the backdoor adjustment formula (Figure 2c) and introduce the three-step process of counterfactual generation (Figrue 2d).
@@
To compute such queries, taking our case as an example (Figure 2d), SCMs follow a principled three-step procedure (Pearl and others, 2000):
@@ B.2 Diffusion Models and Conditional Diffusion Models
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are a class of probabilistic generative models that learn to generate data through a two-phase procedure: a forward diffusion process that progressively adds noise to the data, followed by a reverse denoising process that learns to recover the original data distribution. In recent years, diffusion models have become a powerful paradigm for time series modeling, achieving strong performance across a wide range of tasks (Tashiro et al., 2021; Wen et al., 2023; Yuan and Qiao, 2024; Yang et al., 2024; Lin et al., 2024). We begin by introducing the unconditional setting, detailing both the forward and reverse processes. We then introduce the extension of this framework to the conditional setting, where external factors are incorporated to guide the generation process.
@@ B.2.1 Diffusion Models
Diffusion models can also be interpreted as a special class of score-based generative models (Song and Ermon, 2019), which aim to learn the gradient of the log-density of the data at each timestep. This gradient, known as the score function, indicates the direction in which the noisy input $x_{t}$ should be adjusted to increase its likelihood under the marginal distribution $p_{t}(x_{t})$. The score function is formally defined as:
@@ B.2.2 Conditional Diffusion Models
Ho and Salimans (2022) propose Classifier-Free Guidance (CFG), which can be interpreted as a reparameterization of classifier guidance by manipulating the gradient expression of the conditional log-probability. Starting from Bayes’ rule, we can write the second term in Eq 22 as:
@@ C.1 Motivation for a Diffusion Backbone
Our choice of a diffusion backbone is driven by the goals of Int-TSG and CF-TSG. First, under the score-based view, diffusion models learn and approximate scores $s(x,c)=\nabla_{x}\log p(x\mid c)$ (Eq. 1). Given identifiability of the interventional target $p(x\mid\mathrm{do}(c))$ instead of $p(x\mid c)$, our objective naturally extends to the interventional score $s(x,\mathrm{do}(c))=\nabla_{x}\log p(x\mid\mathrm{do}(c))$, which can be derived from the backdoor-adjusted formulation (Eq. 2). The derivation will be introduced in following sections. Second, diffusion models have become a powerful paradigm for time series modeling, achieving strong performance across imputation, generation, and forecasting tasks (Tashiro et al., 2021; Wen et al., 2023; Yuan and Qiao, 2024; Yang et al., 2024; Lin et al., 2024), offering stable training and good coverage compared to likelihood-free adversarial alternatives. These properties make diffusion a pragmatic and principled backbone for our causal TSG agenda.
@@ C.2 Derivation of backdoor-adjusted score function
The summation over latent confounder $e$ is inside the logarithm, preventing direct application of the gradient operator. This is a well-known difficulty in score-based inference under marginalization. To address this, we adopt a classical result often referred to as the log-derivative of a marginal or softmax trick (Mnih and Gregor, 2014), given by:
@@ D.1 Training objective & optimization
In addition, to encourage diversity among environment embeddings, we initialize a learnable environment bank $\mathbf{E}$ with orthogonal initialization, followed by L2 normalization along the feature dimension, and we regularize the learnable environment bank $\mathbf{E}$ with an orthogonality loss:
@@ E.3 Empirical Evidence of Spurious Correlations
Real-world datasets. For the Air Quality dataset (Figure 11a and 11b), the conditional relation with temperature shows moderate between–split shifts, and wind direction exhibits pronounced nonmonotonic drift, both evident in the LOWESS curves and confirmed by the binned means. For the Traffic dataset (Figure 11c and 11d), the association with cloud coverage shows mild shifts, whereas the relation with hourly rainfall is more strongly nonstationary across splits. Note that the missing bins in Figure 11d arise because rain_1h is zero–inflated with a heavy–tailed positive part; under equal–width binning many mid–range intervals contain no observations and are omitted, producing the visible gap.
@@ H.6 Comparison with Related Works
Prior causal generative models are mostly developed for static modalities, e.g., images via label-graph–consistent GANs or modular recombinations (Kocaoglu et al., 2018; Sauer and Geiger, 2021), text via interventions on latent narrative factors (Hu and Li, 2021), and knowledge graphs via hypothetical triple generation (Liu et al., 2021). our work, to the best of our knowledge, is the first attempt to propose and instantiate causal time series generation, where causal queries target entire sequential data under interventions and regime shifts. Methodologically, within an SCM-based formulation, we realize causality through a back-door–adjusted score guidance that injects confounder adjustment directly into the diffusion sampling field, and we couple it with a learnable environment bank to model latent environments that drive nonstationarity. This is different from designs that either intervene on observed, single-shot factors or require acyclicity-constrained latent DAGs as in CausalVAE (Yang et al., 2021).